# ref.bib UUITP-22/24

### v1

@@ A.2 Useful properties
Notice that if we change the variables as $y=\frac{w}{1-w}$, we can can show that this integral is a special case of the that in (A.3), and we obtain
@@
Appendix B Monodromy of $F_{D}^{(n)}$
@@
$\mathcal{H}^{D}(t,c,\xi\epsilon)=[\mathcal{H}^{D}]_{-2}\,\xi^{-2}+[\mathcal{H}^{D}]_{-1}\,\xi^{-1}+[\mathcal{H}^{D}]_{0}\,\xi^{0}+O(\xi)$ (B.1)
@@
$\Delta_{c}\,\mathcal{H}^{D}(t,x,\xi\epsilon)=\sum_{d=-2}^{\infty}\Delta_{c}\,[\mathcal{H}^{D}(t,c,\epsilon)]_{d}\,\xi^{d}\,.$ (B.2)
@@
B.1 Non-equivariant expansion of the quantum Lebesgue measure
@@
$\mathcal{H}^{D}(t,c,\epsilon)=$ $\frac{\mathrm{e}^{-t\cdot M\cdot\epsilon}\,x^{A_{1}\cdot\epsilon}}{H(x,0,z)^{\sum_{i}\epsilon_{i}}}\times\frac{\operatorname{\Gamma}(\sum_{i}\epsilon_{i})\operatorname{\Gamma}(A_{2}\cdot\epsilon)\operatorname{\Gamma}(k\sum_{i}\epsilon_{i}-A_{2}\cdot\epsilon)}{\operatorname{\Gamma}(k\sum_{i}\epsilon_{i})}$ (B.3)
×$\times F^{(k)}_{D}\Big{(}A_{2}\cdot\epsilon,\sum_{i}\epsilon_{1},\dots,\sum_{i}\epsilon_{i},k\sum_{i}\epsilon_{i};1+y_{1}^{-1},\dots,1+y_{k}^{-1}\Big{)}\,.$
@@
$\left[\frac{\mathrm{e}^{-t\cdot M\cdot\epsilon}\,x^{A_{1}\cdot\epsilon}}{H(x,0,z)^{\sum_{i}\epsilon_{i}}}\right]_{n}=\frac{(-1)^{n}}{n!}\Big{(}t\cdot M\cdot\epsilon+c\,A_{1}\cdot\epsilon+\log H(x,0,z)\sum_{i}\epsilon_{i}\Big{)}^{n}$ (B.4)
@@
$\frac{\operatorname{\Gamma}(\sum_{i}\epsilon_{i})\operatorname{\Gamma}(A_{2}\cdot\epsilon)\operatorname{\Gamma}(k\sum_{i}\epsilon_{i}-A_{2}\cdot\epsilon)}{\operatorname{\Gamma}(k\sum_{i}\epsilon_{i})}=\underbrace{\frac{k}{A_{2}\cdot\epsilon\left(k\sum_{i}\epsilon_{i}-A_{2}\cdot\epsilon\right)}}_{O(\epsilon^{-2})}\;\underbrace{-\frac{\gamma k\sum_{i}\epsilon_{i}}{A_{2}\cdot\epsilon\left(k\sum_{i}\epsilon_{i}-A_{2}\cdot\epsilon\right)}}_{O(\epsilon^{-1})}\\ +\underbrace{\frac{k}{12}\left(\frac{(6\gamma^{2}+\pi^{2})\left(\sum_{i}\epsilon_{i}\right)^{2}}{A_{2}\cdot\epsilon\left(k\sum_{i}\epsilon_{i}-A_{2}\cdot\epsilon\right)}-2\pi^{2}\right)}_{O(\epsilon^{0})}+O(\epsilon)\,.$ (B.5)
@@
$(x)_{n}=x(x+1)\cdots(x+i-1)=x\operatorname{\Gamma}(n)\left(1+xH_{n-1}+O(x^{2})\right)\,,\hskip 30.0ptn\geq 1$ (B.6)
@@
$\left[F^{(k)}_{D}\right]_{1}=\sum_{j=1}^{k}\sum_{i_{j}=1}^{\infty}\left[\frac{(a)_{i_{j}}(b_{j})_{i_{j}}}{(c)_{i_{j}}\,i_{j}!}\right]_{1}x_{j}^{i_{j}}=\sum_{j=1}^{k}\frac{ab_{j}}{c}\sum_{i_{j}=1}^{\infty}\frac{x_{j}^{i_{j}}}{i_{j}}=-\frac{a}{c}\sum_{j=1}^{k}b_{j}\log(1-x_{j})\,.$ (B.9)
@@
$\left[F^{(k)}_{D}\right]_{2}$ $=\sum_{j=1}^{k}\sum_{i_{j}=1}^{\infty}\left[\frac{(a)_{i_{j}}(b_{j})_{i_{j}}}{(c)_{i_{j}}\,{i_{j}}!}\right]_{2}x_{j}^{i_{j}}+\sum_{1\leq j_{1}<j_{2}\leq k}\sum_{i_{j_{1}},i_{j_{2}}=1}^{\infty}\left[\frac{(a)_{i_{j_{1}}+i_{j_{2}}}(b_{j_{1}})_{i_{j_{1}}}(b_{j_{2}})_{i_{j_{2}}}}{(c)_{i_{j_{1}}+i_{j_{2}}}\,i_{j_{1}}!i_{j_{2}}!}\right]_{2}x_{j_{1}}^{i_{j_{1}}}x_{j_{2}}^{i_{j_{2}}}$ (B.10)
@@
B.2 Monodromy of the quantum Lebesgue measure

### v2

@@ A.2 Useful properties
Notice that if we change the variables as $y=\frac{w}{1-w}$, we can show that this integral is a special case of the that in (A.3), and we obtain
@@
Appendix B Quantum volumes as equivariant $B$-model integrals
In this Appendix, we collect certain expressions for the three different types of quantum equivariant volumes that play a central role in our work, namely $\mathcal{F}^{D},\mathcal{H}^{D}$ and $\mathcal{K}^{D}$. We show that all of these admit expressions that are strongly reminiscent of $B$-model integrals, by manipulations that parallel those of the derivation of Hori–Vafa mirror curves of toric Calabi–Yau threefolds. This observation points to a mirror interpretation of symplectic cuts, where each of the quantum volumes involved maps to a certain integral in the mirror geometry. More details will be provided below, we postpone a physical interpretation of these identities to a separate publication.
B.1 Symplectic quotient operators
Let us first recall the Euler integral representation
$\operatorname{\Gamma}(z)=\int_{0}^{\infty}\mathrm{e}^{-u}u^{z-1}\mathop{}\!\mathrm{d}u\,.$ (B.1)
$f(\widehat{u})\cdot\operatorname{\Gamma}(z)=\int_{0}^{\infty}\mathrm{e}^{-u}u^{z-1}f(u)\mathop{}\!\mathrm{d}u\,,$ (B.2)
for any algebraic function $f(u)$.
$\mathcal{G}_{q}(t)\cdot\Psi(\epsilon_{1},\dots,\epsilon_{n}):=$ $\,\delta\Big{(}t+\sum_{i=1}^{n}q_{i}\frac{\partial}{\partial\epsilon_{i}}\Big{)}\cdot\Psi(\epsilon_{1},\dots,\epsilon_{n})$ (B.3)
$\mathcal{G}_{q}(t)\cdot\prod_{i}\operatorname{\Gamma}(\epsilon_{i})=\delta\Big{(}\log\Big{(}\mathrm{e}^{t}\prod_{i}\widehat{u}_{i}{}^{q_{i}}\Big{)}\Big{)}\cdot\prod_{i}\operatorname{\Gamma}(\epsilon_{i})$ (B.4)
$\mathcal{G}_{q}(t)^{-1}\cdot f(t,\epsilon):=\int_{\mathbb{R}}\mathop{}\!\mathrm{d}t\,f(t,\epsilon)\,.$ (B.5)
B.2 Relations among quantum equivariant volumes
$\mathcal{F}^{D}(t,\epsilon)=\Big{(}\prod_{a=1}^{r}\mathcal{G}_{Q^{a}}(t^{a})\Big{)}\cdot\prod_{i=1}^{r+3}\operatorname{\Gamma}(\epsilon_{i})$ (B.6)
$\mathcal{H}^{D}(t,c_{\alpha},\epsilon)=\mathcal{G}_{q^{\alpha}}(c_{\alpha})\cdot\mathcal{F}^{D}(t,\epsilon)\,.$ (B.7)
$\mathcal{K}^{D}(t,c_{1},c_{2},\epsilon):=\Big{(}\prod_{\alpha=1,2}\mathcal{G}_{q^{\alpha}}(c_{\alpha})\Big{)}\cdot\mathcal{F}^{D}(t,\epsilon)$ (B.8)
These relations can be inverted by integrating over the moment map moduli using the operator (B.5). We find
$\int_{\mathbb{R}}\mathop{}\!\mathrm{d}c_{2}\,\mathcal{K}^{D}(t,c_{1},c_{2},\epsilon)$ $=\mathcal{H}^{D}(t,c_{1},\epsilon)$ (B.9)
B.3 Equivariant $B$-model integral for $\mathcal{K}^{D}$
We now give explicit integral formulae for the quantum equivariant volumes of the CY3 and of its cuts. These fit into the hierarchy given in (B.9), with the most fundamental object being $\mathcal{K}^{D}$, which will be out starting point.
We begin by rewriting (4.11) by means of the integral representation (B.1) of Euler’s $\operatorname{\Gamma}$ function and in terms of symplectic quotient operators (B.3)
$\mathcal{K}^{D}(t,c_{1},c_{2},\epsilon)=\prod_{\alpha=1,2}\int_{\mathrm{i}\mathbb{R}}\frac{\mathop{}\!\mathrm{d}\varphi_{\alpha}}{2\pi\mathrm{i}}\mathrm{e}^{c_{\alpha}\varphi_{\alpha}}\prod_{a=1}^{r}\int_{\mathrm{i}\mathbb{R}}\frac{\mathop{}\!\mathrm{d}\phi_{a}}{2\pi\mathrm{i}}\mathrm{e}^{t^{a}\phi_{a}}\prod_{i=1}^{r+3}\int_{0}^{\infty}\frac{\mathop{}\!\mathrm{d}u_{i}}{u_{i}}\,u_{i}^{\epsilon_{i}+\sum_{a}Q^{a}_{i}\,\phi_{a}+\sum_{\alpha}q^{\alpha}_{i}\,\varphi_{\alpha}}\,\mathrm{e}^{-u_{i}}\\ =\prod_{i=1}^{r+3}\int_{0}^{\infty}\frac{\mathop{}\!\mathrm{d}u_{i}}{u_{i}}\,u_{i}^{\epsilon_{i}}\,\mathrm{e}^{-u_{i}}\,\prod_{a=1}^{r}\delta\Big{(}t^{a}+\sum_{i}Q^{a}_{i}\log u_{i}\Big{)}\prod_{\alpha=1,2}\delta\Big{(}c_{\alpha}+\sum_{i}q^{\alpha}_{i}\log u_{i}\Big{)}\,.$ (B.10)
$\mathcal{K}^{D}(\widetilde{t},\epsilon)=\prod_{i=1}^{r+3}\int_{0}^{\infty}\frac{\mathop{}\!\mathrm{d}u_{i}}{u_{i}}\,u_{i}^{\epsilon_{i}}\,\mathrm{e}^{-u_{i}}\,\prod_{a=1}^{r+2}\delta\Big{(}\widetilde{t}^{a}+\sum_{i}\widetilde{Q}^{a}_{i}\log u_{i}\Big{)}\,.$ (B.11)
$\log u_{i}=-\sum_{a=1}^{r+2}\left[\widetilde{Q}(j)^{-1}\right]^{i}_{a}(\widetilde{t}^{a}+\widetilde{Q}^{a}_{j}\log u_{j})\hskip 30.0pt\forall i\neq j\,,$ (B.12)
$u_{i}=\prod_{a=1}^{r+2}\left(\widetilde{z}_{a}^{-1}u_{j}^{\widetilde{Q}^{a}_{j}}\right)^{-\left[\widetilde{Q}(j)^{-1}\right]^{i}_{a}}\,.$ (B.13)
This allows us to reduce the integral to a single variable
$\mathcal{K}^{D}(\widetilde{t},\epsilon)=\int_{0}^{\infty}\,\frac{\mathop{}\!\mathrm{d}u_{j}}{u_{j}^{1-\epsilon_{j}}}\left(\prod_{a=1}^{r+2}(\widetilde{z}_{a}^{-1}u_{j}^{\widetilde{Q}^{a}_{j}})^{-\sum_{i\neq j}\left[\widetilde{Q}(j)^{-1}\right]^{i}_{a}\epsilon_{i}}\right)\mathrm{e}^{-u_{j}-\sum_{i\neq j}\prod_{a=1}^{r+2}(\widetilde{z}_{a}^{-1}u_{j}^{\widetilde{Q}^{a}_{j}})^{-\left[\widetilde{Q}(j)^{-1}\right]^{i}_{a}}}\\ =\mathrm{e}^{-\sum_{a=1}^{r+2}\sum_{i\neq j}\widetilde{t}^{a}\left[\widetilde{Q}(j)^{-1}\right]^{i}_{a}\epsilon_{i}}\int_{0}^{\infty}\frac{\mathop{}\!\mathrm{d}u_{j}}{u_{j}}\,u_{j}^{\epsilon_{j}-\sum_{a=1}^{r+2}\sum_{i\neq j}\widetilde{Q}^{a}_{j}\left[\widetilde{Q}(j)^{-1}\right]^{i}_{a}\epsilon_{i}}\mathrm{e}^{-u_{j}-\sum_{i\neq j}\prod_{a=1}^{r+2}(\widetilde{z}_{a}^{-1}u_{j}^{\widetilde{Q}^{a}_{j}})^{-\left[\widetilde{Q}(j)^{-1}\right]^{i}_{a}}}\,.$ (B.14)
  • $[R(j)]^{i}_{k}=\delta^{i}_{k}$, where $R(j)$ is the square matrix obtained from $R$ by removing the $j$-th column. This trivially follows by noting that $R(j)=\widetilde{Q}(j)^{-1}\widetilde{Q}(j)$.
    $R^{i}_{j}=\sum_{a=1}^{r+2}\left[\widetilde{Q}(j)^{-1}\right]^{i}_{a}\widetilde{Q}_{j}^{a}=-\sum_{k\neq j}\sum_{a=1}^{r+2}\left[\widetilde{Q}(j)^{-1}\right]^{i}_{a}\widetilde{Q}_{k}^{a}=-\sum_{k\neq j}\delta^{i}_{k}=-1$ (B.15)
Using these facts, the above expression simplifies to
$\mathcal{K}^{D}(\widetilde{t},\epsilon)$ $=\mathrm{e}^{-\sum_{a=1}^{r+2}\sum_{i\neq j}\widetilde{t}^{a}\left[\widetilde{Q}(j)^{-1}\right]^{i}_{a}\epsilon_{i}}\int_{0}^{\infty}\frac{\mathop{}\!\mathrm{d}u_{j}}{u_{j}}\,u_{j}^{\sum_{i=1}^{r+3}\epsilon_{i}}\,\exp\left(-u_{j}\left(1+\sum_{i\neq j}\prod_{a=1}^{r+2}\widetilde{z}_{a}^{\left[\widetilde{Q}(j)^{-1}\right]^{i}_{a}}\right)\right)$ (B.16)
Let us define
$H(x_{1},x_{2},z):=1+\sum_{i\neq j}\prod_{a=1}^{r+2}\widetilde{z}_{a}^{\left[\widetilde{Q}(j)^{-1}\right]^{i}_{a}}$ (B.17)
and
$A_{\alpha}^{i}:=\left\{\begin{array}[]{ll}\left[\widetilde{Q}(j)^{-1}\right]^{i}_{r+\alpha}&\text{if $i\neq j$}\\ 0&\text{if $i=j$}\end{array}\right.\,,\hskip 30.0ptM_{a}^{i}:=\left\{\begin{array}[]{ll}\left[\widetilde{Q}(j)^{-1}\right]^{i}_{a}&\text{if $i\neq j$, $a=1,\dots,r$}\\ 0&\text{if $i=j$}\end{array}\right.$ (B.18)
$\mathcal{K}^{D}(t,c_{1},c_{2},\epsilon)$ $=\mathrm{e}^{-t\cdot M\cdot\epsilon}x_{1}^{A_{1}\cdot\epsilon}x_{2}^{A_{2}\cdot\epsilon}\int_{0}^{\infty}\frac{\mathop{}\!\mathrm{d}u_{j}}{u_{j}}\,u_{j}^{\sum_{i=1}^{r+3}\epsilon_{i}}\mathrm{e}^{-u_{j}H(x_{1},x_{2},z)}$ (B.19)
B.3.1 Example: double cut of $\mathbb{C}^{3}$
In the case of a double cut of $\mathbb{C}^{3}$ defined as in Section 4.5.1, the symplectic quotient $X_{1}$ is associated to the charge matrix
$\widetilde{Q}=\left[\begin{array}[]{ccc}0&1&-1\\ 1&-1&0\\ \end{array}\right]$ (B.20)
which corresponds to the matrix obtained after removing the last four columns from (4.31). This precisely corresponds to the fact that $X_{1}$ is the intersection of the last four toric divisors inside of $X_{5}$.
The computation of $\mathcal{K}^{D}$ as a $B$-model integral requires us to make a choice of how to solve the moment map constraints. We choose to solve w.r.t. the first two variables $u_{1}$ and $u_{2}$ in (B.11) as functions of $u_{3}$. As discussed in the previous section, this is equivalent to the choice of index $j=3$ and reduced matrix
$\widetilde{Q}(3)=\left[\begin{array}[]{cc}0&1\\ 1&-1\\ \end{array}\right].$ (B.21)
Substituting into (B.17) and (B.18), we obtain the identifications
$A_{1}\cdot\epsilon=\epsilon_{1}+\epsilon_{2}\,,\hskip 30.0ptA_{2}\cdot\epsilon=\epsilon_{1}\,,\hskip 30.0ptM=0\hskip 30.0pt\text{and}\hskip 30.0ptH(x_{1},x_{2})=1+x_{1}+x_{1}x_{2}\,.$ (B.22)
Finally, the quantum volume of $X_{1}$ can be computed via (B.19), which gives
$\mathcal{K}^{D}(c_{1},c_{2},\epsilon)$ $=\operatorname{\Gamma}(\epsilon_{1}+\epsilon_{2}+\epsilon_{3})\frac{x_{1}^{\epsilon_{1}+\epsilon_{2}}x_{2}^{\epsilon_{1}}}{(1+x_{1}(1+x_{2}))^{\epsilon_{1}+\epsilon_{2}+\epsilon_{3}}}\,.$ (B.23)
B.3.2 Example: double cut of local $\mathbb{P}^{2}$
The double cut of local $\mathbb{P}^{2}$ chosen as in Section 4.5.2 leads to a CY onefold $X_{1}$ with charge matrix
$\widetilde{Q}=\left[\begin{array}[]{cccc|c}z_{1}&z_{2}&z_{3}&z_{4}\\ \hline\cr 1&1&1&-3&t\\ 0&0&-1&1&c_{1}\\ 0&1&-1&0&c_{2}\\ \end{array}\right]\,.$ (B.24)
Choosing to solve the moment map constraints w.r.t. the first three variables $u_{i}$ as functions of $u_{4}$, we obtain the following identifications of parameters:
$A_{1}\cdot\epsilon=2\epsilon_{1}-\epsilon_{2}-\epsilon_{3}\,,\hskip 30.0ptA_{2}\cdot\epsilon=-\epsilon_{1}+\epsilon_{2}\,,\hskip 30.0ptt\cdot M\cdot\epsilon=t\epsilon_{1}$ (B.25)
and
$H(x_{1},x_{2},z)=1+x_{1}^{-1}+zx_{1}^{2}x_{2}^{-1}+x_{1}^{-1}x_{2}\,.$ (B.26)
Plugging in (B.19) we obtain the quantum volume of $X_{1}$.
B.4 $H(x_{1},x_{2})$ vs the Hori–Vafa mirror curve
We recall the standard definition of the Hori–Vafa mirror curve of a toric CY threefold. Let $X_{3}$ be a toric quotient as in (2.1), then the mirror geometry can be described in terms of $r+3$ complex coordinates $Y^{i}\in\mathbb{C}^{\times}$ subject to the constraints
$\sum_{i=1}^{r+3}Q_{i}^{a}Y^{i}=-t^{a}\,,\hskip 30.0pta=1,\dots,r\,,$ (B.27)
whose solutions form a three-dimensional family (see [Katz:1996fh, Hori:2000kt]). Because of the CY condition on $Q_{i}^{a}$, the space of solution is invariant under a simultaneous translation of all the coordinates as
$Y^{i}\mapsto Y^{i}+s$ (B.28)
therefore we can quotient by this action to reduce the space of solutions to a two-dimensional family parametrized by two coordinates $x,y\in\mathbb{C}^{\times}$. Solving explicitly the constraints w.r.t. the $Y^{i}\cong Y^{i}(x,y)$ as functions of $x$ and $y$, we can define the mirror CY manifold as the complex hypersurface
$UV=\sum_{i=1}^{r+3}\mathrm{e}^{Y^{i}(x,y)}=:F(x,y)\,,$ (B.29)
where $U,V\in\mathbb{C}$ are auxiliary complex variables. Because $F(x,y)$ depends on an explicit choice of coordinates for the two-dimensional family of solutions, we have an ambiguity in defining this hypersurface, related to the underlying freedom in the way we parametrize the variables $Y^{i}$ as functions of $x$ and $y$. This ambiguity is of the same nature as the one observed in solving the Dirac-delta constraints in (B.11). In fact, we can formally identify the variables $Y^{i}=\log u_{i}$, where $u_{i}$ are the integration variables in a $B$-model integral representation of the quantum volume of $X_{3}$,
$\mathcal{F}^{D}(t,\epsilon)=\prod_{i=1}^{r+3}\int_{0}^{\infty}\frac{\mathop{}\!\mathrm{d}u_{i}}{u_{i}}\,u_{i}^{\epsilon_{i}}\,\mathrm{e}^{-u_{i}}\,\prod_{a=1}^{r}\delta\Big{(}t^{a}+\sum_{i}Q^{a}_{i}\log u_{i}\Big{)}\,,$ (B.30)
so that the constraints in (B.27) match with the moment map equations defining $X_{3}$.
Similarly, a Lagrangian toric brane $L$ in $X_{3}$ is mapped under mirror symmetry [Hori:2000ic, Aganagic:2001nx] to a mirror brane given by the equation
$U=0=F(x,y)\,,$ (B.31)
which defines a holomorphic submanifold of the hypersurface (B.29). Its moduli space is complex one-dimensional and it corresponds to the curve
$F(x,y)=0$ (B.32)
which is known as the mirror curve. In general, the open string modulus $c$ of the brane can be identified with a function of the two coordinates $x,y$, so that both $x$ and $y$ are completely fixed by these two equations. In practice, the mirror curve is defined by solving the moment map constraints in such a way that $x$ can be directly identified with $\mathrm{e}^{-c}$, while $y=y(x)$ becomes implicitly defined as a solution to $F(x,y)=0$. The specific data associated to the location and orientation of the toric brane $L$ are then encoded by the explicit form of the function $F(x,y)$.
The polynomial $H(x_{1},x_{2})$ appearing in the formula for $\mathcal{K}^{D}$ is related to, but does not quite coincide with, the Hori–Vafa mirror curve $F(x,y)$. More precisely, since $\mathcal{K}^{D}$ is defined by a double cut, which corresponds a pair of toric Lagrangians, there are two mirror curves that describe the moduli spaces of branes $L_{1}$ and $L_{2}$ associated to the cut.
First of all, the branes may come with independent framings and different choices of coordinates. Therefore their moduli spaces would be given by different expressions $F_{1}(x_{1},y_{1})=0$ and $F_{2}(x_{2},y_{2})=0$, with $x_{\alpha}=\mathrm{e}^{-c_{\alpha}}$, $\alpha=1,2$. Neither of these functions is equal to $H(x_{1},x_{2})$ on general grounds. However, all three polynomials are related, since they all come from resolving the moment map constraints, on the universal curve $\sum_{i=1}^{r+3}\mathrm{e}^{Y^{i}}=0$. We illustrate this point with an explicit computation.
An example.
$\left[\begin{array}[]{ccc|c}z_{1}&z_{2}&z_{3}&\\ \hline\cr 0&1&-1&c_{1}\\ 1&-1&0&c_{2}\\ \end{array}\right]$ (B.33)
$u_{i}=\exp(-p_{i}+\mathrm{i}\theta_{i})\,,\hskip 30.0pti=1,2,3$ (B.34)
and use the moment map equations
$p_{2}-p_{3}=c_{1}\,,\hskip 30.0ptp_{1}-p_{2}=c_{2}\,,$ (B.35)
to identify the open string moduli as follows
$x_{1}:=\mathrm{e}^{-c_{1}}=\frac{u_{2}}{u_{3}}\,,\hskip 30.0ptx_{2}:=\mathrm{e}^{-c_{2}}=\frac{u_{1}}{u_{2}}\,.$ (B.36)
$y_{1}=\frac{u_{1}}{u_{3}}\,,\hskip 30.0pty_{2}=\frac{u_{3}}{u_{2}}\,.$ (B.37)
The mirror curves for each brane are obtained by Hori–Vafa as follows
$\begin{split}u_{1}+u_{2}+u_{3}&=u_{3}(1+y_{1}+x_{1})=:u_{3}F_{1}(x_{1},y_{1})\\ &=u_{2}(1+y_{2}+x_{2})=:u_{2}F_{2}(x_{2},y_{2})\\ \end{split}$ (B.38)
$x_{2}=y_{1}x_{1}^{-1}\,,\hskip 30.0pty_{2}=x_{1}^{-1}$ (B.39)
$\begin{split}&p_{2}-p_{3}=-\log u_{2}+\log u_{3}=c_{1}\quad\Rightarrow\quad u_{2}=u_{3}\,\mathrm{e}^{-c_{1}}=u_{3}x_{2}\\ &p_{1}-p_{2}=-\log u_{1}+\log u_{2}=c_{2}\quad\Rightarrow\quad u_{1}=u_{2}\,\mathrm{e}^{-c_{2}}=u_{3}x_{1}x_{2}\,.\end{split}$ (B.40)
The Hori–Vafa universal curve can be written as
$u_{1}+u_{2}+u_{3}=u_{3}(x_{1}x_{2}+x_{2}+1)=:u_{3}H(x_{1},x_{2})$ (B.41)
$H(x_{1},x_{2})=F_{1}(x_{1},y_{1})=y_{2}^{-1}F_{2}(x_{2},y_{2})\,.$ (B.42)
B.5 Equivariant $B$-model integrals for $\mathcal{H}^{D}$ and $\mathcal{F}^{D}$
From (B.19), we can immediately obtain $\mathcal{H}^{D}$ and $\mathcal{F}^{D}$ by integration (B.9)
$\mathcal{H}^{D}(t,c_{1},\epsilon)=\operatorname{\Gamma}\Big{(}\sum_{i}\epsilon_{i}\Big{)}\,\mathrm{e}^{-t\cdot M\cdot\epsilon}\,x_{1}^{A_{1}\cdot\epsilon}\int_{0}^{\infty}\frac{\mathop{}\!\mathrm{d}x_{2}}{x_{2}}\frac{x_{2}^{A_{2}\cdot\epsilon}}{H(x_{1},x_{2},z)^{\sum_{i}\epsilon_{i}}}\,,$ (B.43)
and
$\mathcal{F}^{D}(t,\epsilon)=\operatorname{\Gamma}\Big{(}\sum_{i}\epsilon_{i}\Big{)}\,\mathrm{e}^{-t\cdot M\cdot\epsilon}\int_{0}^{\infty}\frac{\mathop{}\!\mathrm{d}x_{1}}{x_{1}}\int_{0}^{\infty}\frac{\mathop{}\!\mathrm{d}x_{2}}{x_{2}}\frac{x_{1}^{A_{1}\cdot\epsilon}x_{2}^{A_{2}\cdot\epsilon}}{H(x_{1},x_{2},z)^{\sum_{i}\epsilon_{i}}}\,.$ (B.44)
From the discussion of Appendix B.4, it follows that we can replace $H(x,y,z)$ by the mirror curve $F(x,y)$ of either brane defining the double cut (or of the brane defining the single cut associated to $\mathcal{H}^{D}$). The formula remains the same, except that $A_{\alpha},M$ get shifted by some factors, see (B.42) for an example.
Appendix C Monodromy of $F_{D}^{(n)}$
@@
$\mathcal{H}^{D}(t,c,\xi\epsilon)=[\mathcal{H}^{D}]_{-2}\,\xi^{-2}+[\mathcal{H}^{D}]_{-1}\,\xi^{-1}+[\mathcal{H}^{D}]_{0}\,\xi^{0}+O(\xi)$ (C.1)
@@
$\Delta_{c}\,\mathcal{H}^{D}(t,x,\xi\epsilon)=\sum_{d=-2}^{\infty}\Delta_{c}\,[\mathcal{H}^{D}(t,c,\epsilon)]_{d}\,\xi^{d}\,.$ (C.2)
@@
C.1 Non-equivariant expansion of the quantum Lebesgue measure
@@
$\mathcal{H}^{D}(t,c,\epsilon)=$ $\frac{\mathrm{e}^{-t\cdot M\cdot\epsilon}\,x^{A_{1}\cdot\epsilon}}{H(x,0,z)^{\sum_{i}\epsilon_{i}}}\times\frac{\operatorname{\Gamma}(\sum_{i}\epsilon_{i})\operatorname{\Gamma}(A_{2}\cdot\epsilon)\operatorname{\Gamma}(k\sum_{i}\epsilon_{i}-A_{2}\cdot\epsilon)}{\operatorname{\Gamma}(k\sum_{i}\epsilon_{i})}$ (C.3)
×$\times F^{(k)}_{D}\Big{(}A_{2}\cdot\epsilon,\sum_{i}\epsilon_{i},\dots,\sum_{i}\epsilon_{i},k\sum_{i}\epsilon_{i};1+y_{1}^{-1},\dots,1+y_{k}^{-1}\Big{)}\,.$
@@
$\left[\frac{\mathrm{e}^{-t\cdot M\cdot\epsilon}\,x^{A_{1}\cdot\epsilon}}{H(x,0,z)^{\sum_{i}\epsilon_{i}}}\right]_{n}=\frac{(-1)^{n}}{n!}\Big{(}t\cdot M\cdot\epsilon+c\,A_{1}\cdot\epsilon+\log H(x,0,z)\sum_{i}\epsilon_{i}\Big{)}^{n}$ (C.4)
@@
$\frac{\operatorname{\Gamma}(\sum_{i}\epsilon_{i})\operatorname{\Gamma}(A_{2}\cdot\epsilon)\operatorname{\Gamma}(k\sum_{i}\epsilon_{i}-A_{2}\cdot\epsilon)}{\operatorname{\Gamma}(k\sum_{i}\epsilon_{i})}=\underbrace{\frac{k}{A_{2}\cdot\epsilon\left(k\sum_{i}\epsilon_{i}-A_{2}\cdot\epsilon\right)}}_{O(\epsilon^{-2})}\;\underbrace{-\frac{\gamma k\sum_{i}\epsilon_{i}}{A_{2}\cdot\epsilon\left(k\sum_{i}\epsilon_{i}-A_{2}\cdot\epsilon\right)}}_{O(\epsilon^{-1})}\\ +\underbrace{\frac{k}{12}\left(\frac{(6\gamma^{2}+\pi^{2})\left(\sum_{i}\epsilon_{i}\right)^{2}}{A_{2}\cdot\epsilon\left(k\sum_{i}\epsilon_{i}-A_{2}\cdot\epsilon\right)}-2\pi^{2}\right)}_{O(\epsilon^{0})}+O(\epsilon)\,.$ (C.5)
@@
$(x)_{n}=x(x+1)\cdots(x+i-1)=x\operatorname{\Gamma}(n)\left(1+xH_{n-1}+O(x^{2})\right)\,,\hskip 30.0ptn\geq 1$ (C.6)
@@
$\left[F^{(k)}_{D}\right]_{1}=\sum_{j=1}^{k}\sum_{i_{j}=1}^{\infty}\left[\frac{(a)_{i_{j}}(b_{j})_{i_{j}}}{(c)_{i_{j}}\,i_{j}!}\right]_{1}x_{j}^{i_{j}}=\sum_{j=1}^{k}\frac{ab_{j}}{c}\sum_{i_{j}=1}^{\infty}\frac{x_{j}^{i_{j}}}{i_{j}}=-\frac{a}{c}\sum_{j=1}^{k}b_{j}\log(1-x_{j})\,.$ (C.9)
@@
$\left[F^{(k)}_{D}\right]_{2}$ $=\sum_{j=1}^{k}\sum_{i_{j}=1}^{\infty}\left[\frac{(a)_{i_{j}}(b_{j})_{i_{j}}}{(c)_{i_{j}}\,{i_{j}}!}\right]_{2}x_{j}^{i_{j}}+\sum_{1\leq j_{1}<j_{2}\leq k}\sum_{i_{j_{1}},i_{j_{2}}=1}^{\infty}\left[\frac{(a)_{i_{j_{1}}+i_{j_{2}}}(b_{j_{1}})_{i_{j_{1}}}(b_{j_{2}})_{i_{j_{2}}}}{(c)_{i_{j_{1}}+i_{j_{2}}}\,i_{j_{1}}!i_{j_{2}}!}\right]_{2}x_{j_{1}}^{i_{j_{1}}}x_{j_{2}}^{i_{j_{2}}}$ (C.10)
@@
C.2 Monodromy of the quantum Lebesgue measure